# The Right to be Forgotten in Pruning: Unveil Machine Unlearning on Sparse Models

### v1

@@ Abstract
Machine unlearning aims to efficiently eliminate the memory about deleted data from trained models and address the right to be forgotten. Despite the success of existing unlearning algorithms, unlearning in sparse models has not yet been well studied. In this paper, we empirically find that the deleted data has an impact on the pruned topology in a sparse model. Motivated by the observation and the right to be forgotten, we define a new terminology “un-pruning" to eliminate the impact of deleted data on model pruning. Then we propose an un-pruning algorithm to approximate the pruned topology driven by retained data. We remark that any existing unlearning algorithm can be integrated with the proposed un-pruning workflow and the error of un-pruning is upper-bounded in theory. Also, our un-pruning algorithm can be applied to both structured sparse models and unstructured sparse models. In the experiment, we further find that Membership Inference Attack (MIA) accuracy is unreliable for assessing whether a model has forgotten deleted data, as a small change in the amount of deleted data can produce arbitrary MIA results. Accordingly, we devise new performance metrics for sparse models to evaluate the success of un-pruning. Lastly, we conduct extensive experiments to verify the efficacy of un-pruning with various pruning methods and unlearning algorithms. Our code is released at https://anonymous.4open.science/r/UnlearningSparseModels-FBC5/.
@@ 1 Introduction
The Right to be Forgotten is an emerging principle in artificial intelligence (AI) outlined by regulations such as the General Data Protect Regulation (GDPR), the California Consumer Privacy Act (CCPA), and Canada’s proposed Consumer Privacy Protection Act (CPPA) [1, 2, 3]. A straightforward “forgetting” strategy is to retrain new models from scratch on the remaining data as if the deleted data has never been seen by the model. However, the retraining method is impractical due to the expensive cost of frequent deletion requests over complex models. To resolve this challenge, a plethora of machine unlearning methods have been developed to efficiently update the model without retraining [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21].
@@
Despite the success of existing unlearning algorithms, unlearning on sparse models has not been extensively studied yet. We remark that unlearning on sparse models is nontrivial. In Figure 1(a), we visualize the difference between a sparse model based on the original data and a retrained model based on the remaining data (5% data deletion). Specifically, we leverage the Lottery Ticket Hypothesis (LTH) [22] to prune a model with the original training data and the remaining data respectively. We can observe that most indexes of the pruned parameters are different between two sparse models. There is only around a 58% overlap of pruned indexes across two sparse models (60% sparsity) of a ResNet-18. The observation also holds for structured pruning. As shown in Figure 1(b), there are significant differences regarding pruned channels. The results indicate that the pruned indexes are data-dependent.
@@
Motivated by this observation, we raise a new research problem: how to eliminate the influence of deleted data on the pruned topology in a sparse model?More specifically, the problem is how to generate a new sparse model as if the model has never seen the forgotten data. We define “un-pruning" as a new terminology for eliminating the impact of deleted data on pruning. Table 1 lists the difference between un-pruning and other terminologies such as unlearning. We highlight that un-pruning is non-trivial because: (1) while retraining + repruning serves as the gold standard, it is impractical to perform retraining + repruning on a large-scale model for frequent data deletions. (2) existing unlearning algorithms on the remaining parameters in a sparse model cannot change the pruned topology. These algorithms based on approximating activation values may only align the model’s accuracy with the retrained model but fail to align the topology with the standard retraining + repruning [23, 24]; (3) users have the right to withdraw the influence of their data on pruning processes. Our empirical results demonstrate that deleted data has a significant impact on the topology of the sparse model. Consequently, the continued use of sparse models derived from deleted data instances can be deemed illegal [25]. (4) pruning has been considered as a process of learning from the training data [26, 27, 28]. Correspondingly, un-pruning can be considered as a process of unlearning. Thus, updating the topology of a sparse model provides us with a new way of unlearning. We use  [29] to reconstruct training data with the sparse model. Compared with the original sparse model (ResNet-18, 40% sparsity), the un-pruned model is less accurate in reconstructing the training data.
@@
To verify the effectiveness of un-pruning, we follow existing unlearning works to report TA (test accuracy), UA (unlearning accuracy), and MIA (Membership Inference Attack). However, we empirically find that MIA is not a reliable metric to measure the quality of unlearning. With a small change in the data proportion of the binary classification model (i.e., training data vs. test data), the MIA value can fall anywhere within the range of 0 to 1 randomly. Research in large language models also found that the results of MIA are almost equivalent to guesswork [30, 14, 31, 32]. In this paper, we devise new performance metrics to evaluate the quality of unlearning. Specifically, we calculate the structural similarity between neural network representations [33, 34, 35] to measure the unlearning quality. Since retraining + repruning serves as the golden standard, we calculate the intersection of pruned indexes between a repruned sparse model based on remaining data and an “un-pruned" sparse model at the same sparsity level. Intuitively, a high intersection ratio means a high similarity between two topologies. In the experiment, we verify the effectiveness and efficiency of our algorithms with various unlearning algorithms and various unlearning algorithms. The contribution of this work can be summarized as:
@@ 2 Preliminary
Machine Unlearning. Given a model $A$ trained on a complete dataset $D$, the user can request to remove their data $D_{f}$ from $D$. Machine unlearning aims to eliminate the influence of $D_{f}$ on $A$ and unlearn a new model that forgets what has been learned from $D_{f}$. Denote $D_{r}$ as the remaining data where $D=D_{f}\cup D_{r}$. A naive solution is to retrain a new model $A_{r}$ based on $D_{r}$ Dr) from scratch. However, this solution is impractical due to the high computational cost for frequent data deletions over complex models. Recently, approximate unlearning algorithms [8, 9, 10, 36, 37] have been developed to directly generate a new model $A_{u}$ from $A$ without retraining such that the unlearned model $A_{u}$ approximates $A_{r}$ as though it had never seen $D_{r}$: $A_{r}(\cdot)\approx A_{u}(\cdot)$.
@@
Pruning. Pruning can be categorized into two types: unstructured and structured pruning [38, 39, 40, 41, 42, 43, 44]. Denote $\Theta$ as the parameters of a model $A$. Unstructured pruning aims to associate a binary mask $M$ where $|M|=|\Theta|$ and a zero value in $M$ indicates that the corresponding parameter is masked out by $\Theta\odot M$. For example, the Lottery Ticket Hypothesis (LTH) [22] masks out the parameters with the smallest magnitude. Instead of pruning parameters distributed irregularly on the model framework, structured pruning removes entire filters, channels, or even layers. Given a specific prune ratio and a neural network with $A=\{s_{1},s_{2},\cdots,s_{L}\}$, where $s_{i}$ can be the set of channels, filters, neurons, or dense layers, the target is to search for $A^{\prime}=\{s^{\prime}_{1},s^{\prime}_{2},\cdots,s^{\prime}_{L}\}$ to minimize performance degeneration and maximize speed improvement under the given prune ratio, where $s^{\prime}_{i}\subseteq s^{\prime}_{i}$ and $i\in\{1,2,\cdots,L\}$.
@@
Problem setup. How to eliminate the influence of deleted data on pruning is an open problem. As evidenced by our empirical study in Figure 1, the pruned structure is data-dependent. The user has the right to withdraw the influence of their data on pruning as the continued use of sparsed models pruned based on deleted data instances can be deemed illegal [25]. Let $A$ be the sparse model parametrized by $\Theta\odot M$ where $\Theta$ is the set of weights and $M$ is the mask. Suppose the sparse $A$ is trained and pruned based on the original dataset $D$. Given a sub-dataset $D_{f}\subseteq D$ removed from $D$, the problem is to generate a new sparse model $A_{u}$ parametrized by $\Theta_{u}\odot M$ where $\Theta_{u}$ is the new parameters and $M_{u}$ is the new mask. Considering that the retraining+repruning strategy is impractical for frequent data deletions and expensive pruning strategies such as iterative pruning, our target is to directly generate $\Theta_{u}$ and $M_{u}$ without retraining and repruning such that $M_{u}$ is distinguishable from $M$ and $\Theta_{u}\odot M_{u}$ is indistinguishable from the result of retraining+repruning.
@@ 3 Un-pruning
The first challenge of this research problem is that the pruned parameters are frozen during the training process. Therefore, traditional unlearning algorithms that perform parameter updates can not change the value of frozen parameters. A prior work [45] has also studied unlearning on sparse models. Unfortunately, this work does not update the mask during the unlearning process (i.e., the right to be forgotten in pruning has not been addressed). In this paper, we propose an “un-pruning" algorithm to directly generate the new parameters and the new mask without retraining and repruning. Without loss of generality, Algorithm 1 demonstrates our un-pruning process.
@@
1:  for each iteration t= in $10,1,2,\cdots,T-$ 1 do
@@
In this paper, we have investigated 6 most popular approximate unlearning methods: GradientAscent [46], SCRUB [47], Fisher [48, 49, 50], WoodFisher [51], CertifiedUnlearning [52], SFRON [53]. While these unlearning methods varies on the objective function in the unlearning process, all existing unlearning algorithms can be integrated with our method.
We remark that our un-pruning strategy works for these unstructured turning algorithms that do not require extra parameters for pruning. Similarly, in each iteration, we unfreeze the most important channels, filters, or dense layers. Specifically, we use the same standard of pruning structures to choose important structures. For example, Soft Filter Pruning [54] selects filters to be pruned based on the $l_{2}$ norm of filters in the training stage. Accordingly, we can select the filters to recover based on the $l_{2}$ norm. Note that there are some structured pruning algorithms with extra parameters that are used to learn which part to pruneBiP [55, 56, 57]. For those algorithms with extra learnable parameters, we advocate for new algorithms to address the extra parameters since exiting unlearning algorithms only update the model itself.
Evaluation of un-pruning. While the similarity between the unlearned model and the retrained model is usually used to measure the quality of unlearning algorithms [35], it is still an open problem to measure the structural similarity between the “un-pruned" model and the standard “repruned" model. First, we define the intersection ratio and the union ratio of two pruning models:
@@
Besides the structural similarity, we have also introduced the KL-divergence [48] to measure the difference:
@@
where $(x_{i},y_{i})$ is a data point in $D$ and $l$ is the loss function. Given $k$ components in the model: $\Theta=\Theta_{0},\Theta_{1},...,\Theta_{k}$, we follow Unrolling SGD: [4] to write the update the $j$-th parameter of $i$-th component in pruning iteration $s$ in the unlearning process as:
@@
To estimate the un-pruning error regarding $M$, we approximate the unlearning error in [45] in the un-pruning process. Specifically, we first approximate the loss function using the second-order Taylor expansion around the initial mask $M^{0}=M$:
@@
According to [4], the overall parameters changes can be written as:
@@
To further investigate the KL divergence after un-pruning, we measure the difference between the repruned model and retrained model, we follow [58] to introduce the inequality:
@@
$\text{KL}(Q||P)\leq\frac{1}{2}\sum_{i}(1+k_{i}log(\frac{1+\alpha_{1}^{2}||\Theta_{i}^{t}-\Theta_{i}^{0}||}{k_{i}\sigma_{Q,i}^{2})}),$ (21)
where $k_{i}$ is the number of parameters in module $i$ of $\Theta^{t}$ and $\sigma$ is the standard deviation of Gaussian noise. Then we have the following inequality [59]:
@@ 4 Experiment
In this paper, we have followed  [45] to set up the experiment. Specifically, we have introduced 3 models to be pruned: ResNet-18 and AlexNet, and ViT. The models are trained and pruned on 3 datasets: CIFAR-10, FashionMNIST, and ImageNet. For all the datasets and model architectures, we randomly deleted 10% data. We use (LTH) [22] and Soft Filter Pruning [54] as the default unstructured pruning and structured pruning method respectively. The sparse model can reconstruct training data, suggesting that their reconstruction ability is linked to data memorization [28]. Inspired by this, we evaluate whether a sparse ResNet-18 model (with 40% sparsity) retains training data by testing its reconstruction ability. Besides the intersection between the repruned mask and the “un-pruned" mask, we have also introduced IoM(Intersection over Mask), TA (test accuracy), UA (unlearning accuracy), and MIA. We highlight that MIA is fragile as an evaluation of unlearning. We will demonstrate our findings in Section 5.
@@
Results and analysis. For Table LABEL:tab:reconstruction, Extraction-Score refers to the distance between generated images and real images. SSIM refers to the structural similarity between generated images and real images. The experiments shows that the new (un-pruned) sparse model is less accurate in reconstructing the training data compared with the original sparse model. We integrated various unlearning algorithms including Fisher [48], WoodFisher (WF) [51], Gradient Ascent (GA) [46], Certified Unlearning (CU) [52], SCRUB [47], and SFRON [53] into our un-pruning method. Figure 3,  4 and 5 show the experiment result for 3 sparsity levels (i.e., 40%, 60%, and 95%) with structured pruning and structured pruning, respectively. The x-axis indicates IOU and the y-axis measures UA. Figure 7 visualizes the overall performance including sparsity, TA, and UA. Compared with the original model, the un-pruned model is closer to the retrained + repruned model in terms of both IoM and UA. At the same time, the un-pruned model has a very different topology and UA from the original model. Also, we find that our un-pruning has a better performance on structured pruning. This is because structured topology is easier to approximate than unstructured topology where there is a marginal difference in the parameter’s magnitude within a component (e.g., channel). The results verify that our un-pruning algorithm can generate a new sparse model that is similar to retraining + repruning.
@@
Original initialization vs. random initialization. We analyze experiments with two initialization methods. For the original initialization, we save the original initialization weights and restore the initialization weights in the un-pruning process. For the random initialization, we assign the random values to pruned parameters. As shown in Figure 6, the original initialization outperforms the random initialization. This is because the original initialization provides a more stable parameter distribution, thereby achieving better performance in the IoM metric. The result aligns with the viewpoint of  [60] that for a given initialized network, there exists a supermask that performs best on a specific dataset. The original initialization helps us approach the supertask for the remaining dataset, thereby facilitating the forgetting of the unlearning data.

### v2

@@ Abstract
Machine unlearning aims to efficiently eliminate the memory about deleted data from trained models and address the right to be forgotten. Despite the success of existing unlearning algorithms, unlearning in sparse models has not yet been well studied. In this paper, we empirically find that the deleted data has an impact on the pruned topology in a sparse model. Motivated by the observation and the right to be forgotten, we define a new terminology “un-pruning" to eliminate the impact of deleted data on model pruning. Then we propose an un-pruning algorithm to approximate the pruned topology driven by retained data. We remark that any existing unlearning algorithm can be integrated with the proposed un-pruning workflow and the error of un-pruning is upper-bounded in theory. Also, our un-pruning algorithm can be applied to both structured sparse models and unstructured sparse models. In the experiment, we further find that Membership Inference Attack (MIA) accuracy is unreliable for assessing whether a model has forgotten deleted data, as a small change in the amount of deleted data can produce arbitrary MIA results. Accordingly, we devise new performance metrics for sparse models to evaluate the success of un-pruning. Lastly, we conduct extensive experiments to verify the efficacy of un-pruning with various pruning methods and unlearning algorithms. Our code is released at https://github.com/NKUShaw/SparseModels.
@@ 1 Introduction
The Right to be Forgotten is an emerging principle in artificial intelligence (AI) outlined by regulations such as the General Data Protect Regulation (GDPR), the California Consumer Privacy Act (CCPA), and Canada’s proposed Consumer Privacy Protection Act (CPPA) [biega2021reviving, regulation2016regulation, oag2021ccpa]. A straightforward “forgetting” strategy is to retrain new models from scratch on the remaining data as if the deleted data has never been seen by the model. However, the retraining method is impractical due to the expensive cost of frequent deletion requests over complex models. To resolve this challenge, a plethora of machine unlearning methods have been developed to efficiently update the model without retraining [thudi2022unrolling, schelter2019amnesia, bourtoule2021machine, mahadevan2021certifiable, zhang2024towards, liu2024model, liu2024certified, golatkar2020eternal, xu2023machineunlearningsurvey, pal2025llm, liu2025rethinking, fan2025towards, sun2024forget, zhuang2024uoe, fan2024simplicity, zhou2025decoupled, khalil2025not, spartalis2025lotus].
@@
Despite the success of existing unlearning algorithms, unlearning on sparse models has not been extensively studied yet. We remark that unlearning on sparse models is nontrivial. In Figure 1(a), we visualize the difference between a sparse model based on the original data and a retrained model based on the remaining data (5% data deletion). Specifically, we leverage the Lottery Ticket Hypothesis (LTH) [DBLP:conf/iclr/FrankleC19] to prune a model with the original training data and the remaining data respectively. We can observe that most indexes of the pruned parameters are different between two sparse models. There is only around a 58% overlap of pruned indexes across two sparse models (60% sparsity) of a ResNet-18. The observation also holds for structured pruning. As shown in Figure 1(b), there are significant differences regarding pruned channels. The results indicate that the pruned indexes are data-dependent.
@@
Motivated by this observation, we raise a new research problem: how to eliminate the influence of deleted data on the pruned topology in a sparse model?More specifically, the problem is how to generate a new sparse model as if the model has never seen the forgotten data. We define “un-pruning" as a new terminology for eliminating the impact of deleted data on pruning. Table 1 lists the difference between un-pruning and other terminologies such as unlearning. We highlight that un-pruning is non-trivial because: (1) while retraining + repruning serves as the gold standard, it is impractical to perform retraining + repruning on a large-scale model for frequent data deletions. (2) existing unlearning algorithms on the remaining parameters in a sparse model cannot change the pruned topology. These algorithms based on approximating activation values may only align the model’s accuracy with the retrained model but fail to align the topology with the standard retraining + repruning [wang2024machine, qu2023learn]; (3) users have the right to withdraw the influence of their data on pruning processes. Our empirical results demonstrate that deleted data has a significant impact on the topology of the sparse model. Consequently, the continued use of sparse models derived from deleted data instances can be deemed illegal [DBLP:journals/clsr/VillarongaKL18]. (4) pruning has been considered as a process of learning from the training data [DBLP:conf/nips/ZhouLLY19, DBLP:conf/cvpr/TungM18, haim2022reconstructing]. Correspondingly, un-pruning can be considered as a process of unlearning. Thus, updating the topology of a sparse model provides us with a new way of unlearning. We use  [DBLP:conf/nips/HaimVYSI22] to reconstruct training data with the sparse model. Compared with the original sparse model (ResNet-18, 40% sparsity), the un-pruned model is less accurate in reconstructing the training data.
@@
To verify the effectiveness of un-pruning, we follow existing unlearning works to report TA (test accuracy), UA (unlearning accuracy), and MIA (Membership Inference Attack). However, we empirically find that MIA is not a reliable metric to measure the quality of unlearning. With a small change in the data proportion of the binary classification model (i.e., training data vs. test data), the MIA value can fall anywhere within the range of 0 to 1 randomly. Research in large language models also found that the results of MIA are almost equivalent to guesswork [duan2024membership, liu2025rethinking, das2024blind, zhang2024membership]. In this paper, we devise new performance metrics to evaluate the quality of unlearning. Specifically, we calculate the structural similarity between neural network representations [DBLP:journals/corr/abs-2305-06329, DBLP:conf/icml/Kornblith0LH19, DBLP:journals/csur/XuZZZY24] to measure the unlearning quality. Since retraining + repruning serves as the golden standard, we calculate the intersection of pruned indexes between a repruned sparse model based on remaining data and an “un-pruned" sparse model at the same sparsity level. Intuitively, a high intersection ratio means a high similarity between two topologies. In the experiment, we verify the effectiveness and efficiency of our algorithms with various unlearning algorithms and various unlearning algorithms. The contribution of this work can be summarized as:
@@ 2 Preliminary
Machine Unlearning. Given a model $A$ trained on a complete dataset $D$, the user can request to remove their data $D_{f}$ from $D$. Machine unlearning aims to eliminate the influence of $D_{f}$ on $A$ and unlearn a new model that forgets what has been learned from $D_{f}$. Denote $D_{r}$ as the remaining data where $D=D_{f}\cup D_{r}$. A naive solution is to retrain a new model $A_{r}$ based on $D_{r}$ Dr) from scratch. However, this solution is impractical due to the high computational cost for frequent data deletions over complex models. Recently, approximate unlearning algorithms [zhang2024towards, liu2024model, liu2024certified, DBLP:journals/corr/abs-2407-21035, DBLP:journals/corr/abs-2401-10371] have been developed to directly generate a new model $A_{u}$ from $A$ without retraining such that the unlearned model $A_{u}$ approximates $A_{r}$ as though it had never seen $D_{r}$: $A_{r}(\cdot)\approx A_{u}(\cdot)$.
@@
Pruning. Pruning can be categorized into two types: unstructured and structured pruning [DBLP:journals/pami/ChengZS24a, he2023structured, xu2020convolutional, cheng2024survey, tang2024survey, ma2018survey, deng2020model]. Denote $\Theta$ as the parameters of a model $A$. Unstructured pruning aims to associate a binary mask $M$ where $|M|=|\Theta|$ and a zero value in $M$ indicates that the corresponding parameter is masked out by $\Theta\odot M$. For example, the Lottery Ticket Hypothesis (LTH) [DBLP:conf/iclr/FrankleC19] masks out the parameters with the smallest magnitude. Instead of pruning parameters distributed irregularly on the model framework, structured pruning removes entire filters, channels, or even layers. Given a specific prune ratio and a neural network with $A=\{s_{1},s_{2},\cdots,s_{L}\}$, where $s_{i}$ can be the set of channels, filters, neurons, or dense layers, the target is to search for $A^{\prime}=\{s^{\prime}_{1},s^{\prime}_{2},\cdots,s^{\prime}_{L}\}$ to minimize performance degeneration and maximize speed improvement under the given prune ratio, where $s^{\prime}_{i}\subseteq s^{\prime}_{i}$ and $i\in\{1,2,\cdots,L\}$.
@@
Problem setup. How to eliminate the influence of deleted data on pruning is an open problem. As evidenced by our empirical study in Figure 1, the pruned structure is data-dependent. The user has the right to withdraw the influence of their data on pruning as the continued use of sparsed models pruned based on deleted data instances can be deemed illegal [DBLP:journals/clsr/VillarongaKL18]. Let $A$ be the sparse model parametrized by $\Theta\odot M$ where $\Theta$ is the set of weights and $M$ is the mask. Suppose the sparse $A$ is trained and pruned based on the original dataset $D$. Given a sub-dataset $D_{f}\subseteq D$ removed from $D$, the problem is to generate a new sparse model $A_{u}$ parametrized by $\Theta_{u}\odot M$ where $\Theta_{u}$ is the new parameters and $M_{u}$ is the new mask. Considering that the retraining+repruning strategy is impractical for frequent data deletions and expensive pruning strategies such as iterative pruning, our target is to directly generate $\Theta_{u}$ and $M_{u}$ without retraining and repruning such that $M_{u}$ is distinguishable from $M$ and $\Theta_{u}\odot M_{u}$ is indistinguishable from the result of retraining+repruning.
@@ 3 Un-pruning
The first challenge of this research problem is that the pruned parameters are frozen during the training process. Therefore, traditional unlearning algorithms that perform parameter updates can not change the value of frozen parameters. A prior work [DBLP:conf/nips/JiaLRYLLSL23] has also studied unlearning on sparse models. Unfortunately, this work does not update the mask during the unlearning process (i.e., the right to be forgotten in pruning has not been addressed). In this paper, we propose an “un-pruning" algorithm to directly generate the new parameters and the new mask without retraining and repruning. Without loss of generality, Algorithm 1 demonstrates our un-pruning process.
@@
1:for each iteration t= in $0,1,2,\cdots,T-1$ do
@@
In this paper, we have investigated 6 most popular approximate unlearning methods: GradientAscent [DBLP:conf/aaai/GravesNG21], SCRUB [DBLP:conf/nips/KurmanjiTHT23], Fisher [DBLP:conf/cvpr/GolatkarAS20, foster2024fast, kirkpatrick2017overcoming], WoodFisher [DBLP:conf/nips/SinghA20], CertifiedUnlearning [DBLP:conf/icml/ZhangDWL24], SFRON [huang2024unified]. While these unlearning methods varies on the objective function in the unlearning process, all existing unlearning algorithms can be integrated with our method.
We remark that our un-pruning strategy works for these unstructured turning algorithms that do not require extra parameters for pruning. Similarly, in each iteration, we unfreeze the most important channels, filters, or dense layers. Specifically, we use the same standard of pruning structures to choose important structures. For example, Soft Filter Pruning [he2018soft] selects filters to be pruned based on the $l_{2}$ norm of filters in the training stage. Accordingly, we can select the filters to recover based on the $l_{2}$ norm. Note that there are some structured pruning algorithms with extra parameters that are used to learn which part to pruneBiP [zhang2022advancing, sehwag2020hydra, wang2022trainability]. For those algorithms with extra learnable parameters, we advocate for new algorithms to address the extra parameters since exiting unlearning algorithms only update the model itself.
Evaluation of un-pruning. While the similarity between the unlearned model and the retrained model is usually used to measure the quality of unlearning algorithms [DBLP:journals/csur/XuZZZY24], it is still an open problem to measure the structural similarity between the “un-pruned" model and the standard “repruned" model. First, we define the intersection ratio and the union ratio of two pruning models:
@@
Besides the structural similarity, we have also introduced the KL-divergence [DBLP:conf/cvpr/GolatkarAS20] to measure the difference:
@@
where $(x_{i},y_{i})$ is a data point in $D$ and $l$ is the loss function. Given $k$ components in the model: $\Theta=\Theta_{0},\Theta_{1},...,\Theta_{k}$, we follow Unrolling SGD: [thudi2022unrolling] to write the update the $j$-th parameter of $i$-th component in pruning iteration $s$ in the unlearning process as:
@@
To estimate the un-pruning error regarding $M$, we approximate the unlearning error in [DBLP:conf/nips/JiaLRYLLSL23] in the un-pruning process. Specifically, we first approximate the loss function using the second-order Taylor expansion around the initial mask $M^{0}=M$:
@@
According to [thudi2022unrolling], the overall parameters changes can be written as:
@@
To further investigate the KL divergence after un-pruning, we measure the difference between the repruned model and retrained model, we follow [DBLP:conf/uai/DziugaiteR17] to introduce the inequality:
@@
[…]
where $k_{i}$ is the number of parameters in module $i$ of $\Theta^{t}$ and $\sigma$ is the standard deviation of Gaussian noise. Then we have the following inequality [DBLP:conf/iclr/ChatterjiNS20]:
@@ 4 Experiment
In this paper, we have followed  [DBLP:conf/nips/JiaLRYLLSL23] to set up the experiment. Specifically, we have introduced 3 models to be pruned: ResNet-18 and AlexNet, and ViT. The models are trained and pruned on 3 datasets: CIFAR-10, FashionMNIST, and ImageNet. For all the datasets and model architectures, we randomly deleted 10% data. We use (LTH) [DBLP:conf/iclr/FrankleC19] and Soft Filter Pruning [he2018soft] as the default unstructured pruning and structured pruning method respectively. The sparse model can reconstruct training data, suggesting that their reconstruction ability is linked to data memorization [haim2022reconstructing]. Inspired by this, we evaluate whether a sparse ResNet-18 model (with 40% sparsity) retains training data by testing its reconstruction ability. Besides the intersection between the repruned mask and the “un-pruned" mask, we have also introduced IoM(Intersection over Mask), TA (test accuracy), UA (unlearning accuracy), and MIA. We highlight that MIA is fragile as an evaluation of unlearning. We will demonstrate our findings in Section 5.
@@
Results and analysis. For Table LABEL:tab:reconstruction, Extraction-Score refers to the distance between generated images and real images. SSIM refers to the structural similarity between generated images and real images. The experiments shows that the new (un-pruned) sparse model is less accurate in reconstructing the training data compared with the original sparse model. We integrated various unlearning algorithms including Fisher [DBLP:conf/cvpr/GolatkarAS20], WoodFisher (WF) [DBLP:conf/nips/SinghA20], Gradient Ascent (GA) [DBLP:conf/aaai/GravesNG21], Certified Unlearning (CU) [DBLP:conf/icml/ZhangDWL24], SCRUB [DBLP:conf/nips/KurmanjiTHT23], and SFRON [huang2024unified] into our un-pruning method. Figure 3,  4 and 5 show the experiment result for 3 sparsity levels (i.e., 40%, 60%, and 95%) with structured pruning and structured pruning, respectively. The x-axis indicates IOU and the y-axis measures UA. Figure 7 visualizes the overall performance including sparsity, TA, and UA. Compared with the original model, the un-pruned model is closer to the retrained + repruned model in terms of both IoM and UA. At the same time, the un-pruned model has a very different topology and UA from the original model. Also, we find that our un-pruning has a better performance on structured pruning. This is because structured topology is easier to approximate than unstructured topology where there is a marginal difference in the parameter’s magnitude within a component (e.g., channel). The results verify that our un-pruning algorithm can generate a new sparse model that is similar to retraining + repruning.
@@
Original initialization vs. random initialization. We analyze experiments with two initialization methods. For the original initialization, we save the original initialization weights and restore the initialization weights in the un-pruning process. For the random initialization, we assign the random values to pruned parameters. As shown in Figure 6, the original initialization outperforms the random initialization. This is because the original initialization provides a more stable parameter distribution, thereby achieving better performance in the IoM metric. The result aligns with the viewpoint of  [zhou2019deconstructing] that for a given initialized network, there exists a supermask that performs best on a specific dataset. The original initialization helps us approach the supertask for the remaining dataset, thereby facilitating the forgetting of the unlearning data.